Solve the System by Substitution:
y=x+4
3x+y=16

Answer by rapaljer(4551)

y=x+4
3x+y=16

Substitute the y from the first equation into the second equation:
3x+ (_____)= 16
3x + (x+4) = 16
4x + 4 = 16
4x = 12
x=3

Substitute the x into the "y=" equation in order to find y:
y=x+4
y = 3+ 4= 7

Check in the second equation:
3x+y=16
3(3) + 7= 16
9+7=16
It checks!

R^2 at SCC